OpenStudy (anonymous):

Solve by elimination method x+4y=19 -x+7y=14

OpenStudy (asadkarim7):

eliminate x 11y = 33 y = 33/11 y = 3 substitute y in eq 1 x + 4( 3) = 19 x + 12 = 19 x = 19 -12

OpenStudy (asadkarim7):

x= 7 the solution set ( 7 ,3 )

OpenStudy (anonymous):

Just add 2 equations x+4y-x+7y = 19+14 => 11y = 33 => y = 3 Now from 1st equation x =19 - 4y => x = 19-12 =>x = 7

OpenStudy (anonymous):

-x + 7y = 14 or, -x = 14 - 7y or, x = 7y - 14 Putting the value of x in the first equation, 7y -14 + 4y = 19 or, 11 y = 19 + 14 or, 11y = 33 or, y = 33/11 = 3 so, x = 7*3 -14 = 21 - 14 = 7 So, (x,y) = (7, 3)